Consider the sequence of real numbers 13. 1 1 2 1 2 + 2 + 1 2 + Show that this sequence is convergent and find its limit byProve that any sequence in hence we may suppose that each subsequence has a least term.) (Note that this result and the theorplease do both question clearly,thank you!
Answer #1

& the sequence an= 21 * 2+1 terventimes) Let hun har t = I then bad d = 1 (121) 2 in the sequence is convergent tol ital) (14